Oct 21

Conditions: Thick clouds at sunset, thinner clouds throughout the night,

Observers: Wen-fai Fong, Peter Milne

Instrument specialist: Dave Thompson

01:00 UT: Sunset. Thick clouds, attempt at twilight flats but nothing usable due to clouds

02:00 UT: Start P. Milne's program

04:00 UT: Start W. Fong's program

04:45 UT: Telescope closed due to high humidity (>90%).

07:44 UT: Telescope re-opened. Spent 1st hour trying to focus through clouds and sharp temperature changes.

08:44 UT: Start P. Milne's program.

09:30 UT: Start W. Fong's program.

LBC UI issues:

-"Scale exptimes by" function gives an error, although it still works.

-Sometimes OB upload function doesn't work so you have to re-upload in LBC UI.

-- %USERSIG{EiichiEgami - 2014-10-21}%
Topic revision: r4 - 22 Oct 2014, EiichiEgami
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback